OpenStudy (goldrush18):

ABC Accounting Services provides general accounting, audit and advisory services to clients. ABC charges clients at the rate of $150 per hour for its services. The overhead charge per hour for job is $40. Clients A and B each incurred 120 hours of accounting services during May. Client A's accounting services were conducted by two senior accountants and one junior accountant each working 40 hours. Client B's consulting was done by four junior accountants each working equal time. The actual cost of each class of accountant is Senior Accountant $75 per hour Junior Accountant $47.50 per hour

OpenStudy (goldrush18):

Calculate the total cost of each client job.

OpenStudy (anonymous):

ok. for client A. we have : (2 * senior accountant * working hours)+(1* junior accountant)+120 * 40 = (2 * 75 * 40) + (1 * 47.50 * 40) + 4800 = 6000 + 1900 + 4800 = 12'700 $ income for A is = (120 * 150) - 12700 = 5'300 $

OpenStudy (anonymous):

and for B : (4 * 47.50 * 30) + 120 * 40 = 5700 + 4800 = 10'500 $ income for B is : 120 * 150 - 10500 = 7'500 $

OpenStudy (anonymous):

total cost of each client = paid for accountants + overhead costs

OpenStudy (anonymous):

helo0o0o00o , are you here ?